# Breathing Black Hole Shadows in Modified Gravity (MOG)

### v1

@@ Abstract
In this paper, we investigate the dynamic phenomenological signatures of a Schwarzschild-MOG black hole shadow perturbed by passing gravitational waves. By perturbing the Hamilton-Jacobi equation for photon null geodesics, we demonstrate that the unique field content of MOG breaks the observational degeneracy with standard General Relativity. We mathematically prove two distinct, time-dependent signatures. First, the massless MOG scalar field induces a volumetric “breathing mode” polarization, causing the total apparent area of the shadow to rhythmically expand and contract. Second, the massive MOG vector field undergoes quantum vacuum dispersion, arriving at the observer with a predictable time delay. This delayed massive wave sources secondary longitudinal metric perturbations that manifest as a sudden, asymmetric translational wobble of the shadow on the celestial screen. These dynamic geometric shifts offer a robust observational template for next-generation interferometry to strictly test the existence of massive force carriers and scalar fields in gravity.
@@ I Introduction
Among the most compelling alternative frameworks is Scalar-Tensor-Vector Gravity (STVG), widely referred to as Modified Gravity (MOG) [12]. In MOG, the gravitational interaction is not governed exclusively by a massless tensor field. Instead, the theory introduces an enhanced, dynamical scalar gravitational coupling, alongside a massive vector field that exerts a repulsive fifth force. This theoretical architecture successfully replicates the flat rotation curves of galaxies and the dynamics of galaxy clusters without necessitating cold dark matter [13, 14]. In the strong-field regime, the static MOG black hole solutions yield event horizons and photon spheres whose geometries are distinctly altered by the MOG deformation parameter $\alpha$, producing static observer shadows that differ slightly from their Schwarzschild and Kerr counterparts [15, 16, 17].
While the static shadows of modified gravity theories have been extensively categorized, the universe is inherently dynamic [18]. A fundamental gap remains in understanding how these modified geometries respond when bathed in transient gravitational radiation. When a standard General Relativity gravitational wave washes over a black hole, it perturbs the spacetime metric exclusively through transverse-traceless tensor polarization modes (the “plus” and “cross” modes) [19]. These modes are strictly volume-preserving. Therefore, if a standard gravitational wave dynamically distorts a black hole shadow, it can only stretch and squash the circular boundary into an ellipse, strictly conserving the total apparent area of the shadow on the observer’s celestial screen.
In this paper, we bridge this theoretical gap by mathematically deriving the dynamic interaction between gravitational waves and the shadow of a non-rotating Schwarzschild-MOG black hole. Because MOG introduces extra degrees of freedom into the gravitational sector, its gravitational waves carry novel polarization modes [20, 21]. We utilize the Hamilton-Jacobi formalism for perturbed null geodesics [22, 23] to explicitly map how the MOG scalar and massive vector fields dynamically shift the critical impact parameters of trapped photons. We demonstrate that the scalar field fluctuations generate a strictly volumetric “breathing mode,” meaning the area of the shadow will rhythmically dilate and contract [24, 25], a smoking gun signature forbidden in standard General Relativity.
Furthermore, we explore the phenomenological consequences of the massive vector field. Because massive fields undergo vacuum dispersion [26, 27], they travel at a group velocity strictly less than the speed of light. We show that this dispersion generates a delayed geometric echo. When this massive, delayed wave interacts with the black hole geometry, it sources secondary longitudinal metric perturbations, causing the center of the shadow to experience an asymmetric translational wobble.
@@ II The Unperturbed Schwarzschild-MOG Spacetime
To understand how a gravitational wave dynamically alters a black hole shadow, we must first rigorously define the static, unperturbed background. In this section, we construct the spherically symmetric, non-rotating black hole spacetime in Scalar-Tensor-Vector Gravity (STVG), commonly referred to as Modified Gravity (MOG) [12]. We then determine the exact radius of the photon sphere and the resulting static shadow projected to a distant observer [15].
In standard General Relativity, a non-rotating black hole is described by the Schwarzschild metric. In MOG, the presence of a massive vector field (which exerts a repulsive gravitational force) and an enhanced scalar gravitational constant alters the geometry [12]. The unperturbed Schwarzschild-MOG metric in standard spherical coordinates $(t,r,\theta,\phi)$ is given by the line element:
@@
The boundary of a black hole shadow is defined by the paths of light rays (photons) that barely escape the black hole’s gravity. To trace these paths, we use the Hamilton-Jacobi equation for a photon traveling along a null geodesic (a path where $ds^{2}=0$).
Because the spacetime is spherically symmetric, we can, without loss of generality, confine the photon’s orbit to the equatorial plane by setting $\theta=\pi/2$. This simplifies the metric symmetries, yielding two fundamentally conserved quantities for the photon along its trajectory:
@@
The edge of the black hole shadow is formed by photons trapped in an unstable circular orbit. This region is called the photon sphere. For a photon to remain in a perfect circular orbit at a constant radius $r_{p}$, it must sit precisely at the peak of the effective potential barrier. This requires two simultaneous physical conditions:
@@
Applying the second condition (taking the derivative of Eq. (8) with respect to $r$ and setting it to zero) yields a simple quadratic equation governing the photon sphere radius:
@@
A distant observer does not see $r_{p}$ directly. Instead, they see the apparent shadow projected onto their celestial sky [28]. The size of this dark circular disk is determined by the critical impact parameter, denoted as $b_{c}=L/E$. By rearranging the first condition ($V_{\text{eff}}(r_{p})=E^{2}$), we find that the static shadow radius $R_{\text{sh}}$ is directly tied to the impact parameter:
@@
Substituting our MOG metric function $f(r_{p})$, the final unperturbed shadow radius becomes:
@@
Figure 1: Simultaneous evolution of the black hole shadow radius $R_{\text{sh}}$ (left axis) and the total apparent area fluctuation $\delta A(t)$ (right axis) under the influence of a transient scalar gravitational wave. The navy dashed curves represent the Schwarzschild (GR) baseline, while the crimson and green solid curves depict the enhanced Schwarzschild-MOG response for a deformation parameter $\alpha=0.2$. Both models exhibit a characteristic “breathing mode” triggered at the wave arrival time $t_{0}=5.0$, characterized by an exponentially damped oscillation ($\tau=8.0$) at the quasi-normal mode frequency $\omega_{b}=1.5$. Note that the MOG framework yields a larger static baseline and a higher amplitude of area fluctuation, providing a distinct observational signature for scalar-tensor-vector gravity in the strong-field regime.
In Figure 1, we present the numerical results of the combined dynamics of the shadow radius and area fluctuations. The plot highlights the fundamental coupling between the scalar gravitational strain $h_{b}(t)$ and the geometric observables. By utilizing a dual-axis representation, we demonstrate that the MOG deformation parameter $\alpha$ acts as a scaling factor for the breathing mode; not only is the static shadow larger than the Schwarzschild counterpart, but the magnitude of the dynamic response to the same incident wave is amplified. This synchronization of the radial oscillation and the volumetric area change mathematically derived in Eq. (22) and Eq. (26)—constitutes a primary differentiator for the STVG framework.
@@ III Scalar Gravitational Wave Perturbations: The Breathing Mode
Having established the static Schwarzschild-MOG background, we now introduce a time-dependent perturbation to the spacetime. In standard General Relativity, gravitational waves possess only two transverse-traceless tensor polarization modes, commonly known as the “plus” and “cross” modes. These tensor modes are strictly volume-preserving; if they pass through a circular ring of test particles, they will stretch and squash the circle into an ellipse, but the total area enclosed by the ring remains constant. However, Modified Gravity (MOG) introduces an additional dynamic scalar field, $G$, into the gravitational action [29]. Fluctuations in this scalar field generate a unique scalar gravitational wave polarization known as the “breathing mode” [30, 31]. Unlike tensor modes, the breathing mode is entirely volumetric—it causes space to rhythmically expand and contract uniformly in the transverse plane.
@@
where $\bar{g}_{\mu\nu}$ is our static Schwarzschild-MOG background from Eq. (1), and $h_{\mu\nu}$ is the significantly small perturbation caused by the passing wave. For a pure scalar breathing mode generated by the MOG scalar field, the spatial perturbations in the transverse plane (the $x$-$y$ plane, corresponding to the observer’s screen) take the form of a simple conformal factor. We denote this time-dependent strain as $h_{b}(t)$.
To analytically model the transient scalar strain $h_{b}(t)$, we draw upon the foundational principles of black hole perturbation theory. When a black hole is dynamically perturbed such as during the final stages of a compact binary coalescence the background spacetime geometry acts as an effective potential barrier to incoming wave fields. As first demonstrated by Vishveshwara [32] and generalized by Teukolsky [33], the newly formed or perturbed black hole radiates away these geometric deformations as quasi-normal modes (QNMs). Because the black hole is an open, dissipative system losing energy to both spatial infinity and the event horizon, the characteristic eigenvalues governing these wave equations are strictly complex, taking the form $\tilde{\omega}=\omega_{b}-i/\tau$. Here, the real component $\omega_{b}$ dictates the fundamental oscillation frequency, while the imaginary component yields a characteristic exponential damping time $\tau$.
Translating these complex frequencies into the time domain strictly produces a damped sinusoid. Therefore, to model the incident scalar wave packet as a fundamental QNM ringdown, we define the time-dependent scalar strain $h_{b}(t)$ as a piecewise function. Applying the strict causality constraint of the wave’s arrival time $t_{0}$, the strain is explicitly:
@@
We must now determine how this oscillating metric affects the path of the light rays. The motion of a photon is dictated by the Hamiltonian,
@@
where $p_{\mu}$ is the four-momentum of the photon. We can split this Hamiltonian into two parts: the unperturbed background part ($\bar{H}$) and the small shift caused by the wave ($\delta H$). This is expressed as $H=\bar{H}+\delta H=0$. The unperturbed Hamiltonian yields the static orbits we derived in Section II. The perturbation term, driven entirely by the gravitational wave, is:
@@
Because the wave is traveling along the $z$-axis, it primarily disrupts the angular momentum components of the photon’s momentum, specifically $p_{x}$ and $p_{y}$ in Cartesian coordinates. Substituting our perturbed inverse metric from Eq. (15) into the Hamiltonian perturbation, and recalling that the square of the transverse momentum corresponds to $L^{2}/r^{2}$ in spherical coordinates, we find:
@@
This shift in the Hamiltonian directly modifies the radial equation of motion for the photon. When we substitute the perturbed Hamiltonian back into Eq. (7), the gravitational wave essentially injects a time-dependent fluctuation into the effective potential barrier. The new, total dynamic effective potential is the sum of the unperturbed potential (Equation 8) and the wave perturbation:
@@
This equation reveals a profound physical insight. Because the scalar wave strain $h_{b}(t)$ depends only on time and not on the radial coordinate $r$, the spatial derivative of the effective potential ($\partial_{r}V_{\text{eff}}^{\text{total}}$) is not shifted to first order. Consequently, the physical location of the photon sphere remains perfectly anchored at the static radius $r_{p}$ (Equation 10). The wave washes over the black hole, but the spherical boundary where light is trapped does not change its radial coordinate. However, the energy required for a photon to escape that trapped orbit is dynamically altered by the wave.
To find the new edge of the black hole shadow, we must recalculate the critical impact parameter. We set the peak of the new total effective potential equal to the squared energy of the photon ($E^{2}$) evaluated at the exact photon sphere radius $r_{p}$:
@@
At this juncture, it is crucial to explicitly clarify the physical regime in which this formulation operates. By setting the peak of the time-dependent effective potential equal to the unperturbed squared energy $E^{2}$ in Eq. (20), our derivation implicitly relies on the adiabatic approximation. In a strictly dynamic spacetime where the metric fluctuates with time ($\partial_{t}g_{\mu\nu}\neq 0$), the photon’s total energy $E=-p_{t}$ is not perfectly conserved along its null geodesic. This adiabatic assumption is most robust when the period of the passing gravitational wave is significantly longer than the time it takes for a photon to cross the strong-field region of the photon sphere. We acknowledge that for highly dynamic perturbations such as fundamental quasinormal modes where the wave’s oscillation frequency ($\omega_{b}\sim 1/M$) is comparable to the photon’s orbital frequency the photon can exchange energy with the gravitational wave, meaning $dE/d\lambda\neq 0$. In such extreme regimes, achieving mathematical exactness would require a fully integrated geodesic equation. Nevertheless, evaluating the critical impact parameter through this adiabatic Hamiltonian formalism successfully and elegantly captures the leading-order kinematic deformation of the shadow’s boundary, providing a highly accurate phenomenological baseline. By rearranging this expression to solve for the time-dependent critical impact parameter, $b_{c}(t)=L/E$, we obtain the mathematical formula for the dynamic shadow radius, $R_{\text{sh}}(t)$:
@@
Because the amplitude of a gravitational wave strain is extraordinarily small ($h_{b}\ll 1$), we can safely apply a first-order Taylor expansion to the fraction $(1-h_{b})^{-1}\approx 1+h_{b}(t)$. Taking the square root of both sides, and recognizing that $r_{p}/\sqrt{f(r_{p})}$ is exactly the unperturbed static shadow radius $\bar{R}_{\text{sh}}$ from Equation 12, we arrive at the primary phenomenological result of this paper:
@@
To provide a complete phenomenological template for observational testing, we synthesize the static modified geometry with our dynamic wave model. By substituting the geometric expansion of the unperturbed shadow radius and the quasi-normal mode strain $h_{b}(t)$ into Equation 22, we arrive at the master equation for the dynamic shadow radius:
@@
This expanded expression elegantly decouples the theoretical background from the transient perturbation. The leading square-root term isolates the static Schwarzschild-MOG geometry, explicitly governed by the Newtonian mass $M$, the deformation parameter $\alpha$, and the critical photon sphere $r_{p}$. To fully unpack the localized gravitational wave interaction, it is most instructive to evaluate Equation 23 across its two distinct temporal regimes. Let us explicitly define the unperturbed, static Schwarzschild-MOG shadow radius as the geometric baseline:
@@
In standard General Relativity, a passing gravitational wave perturbs the spacetime metric exclusively through transverse-traceless tensor modes. Consequently, such a wave can only dynamically distort the black hole shadow into an ellipse, strictly conserving its total apparent area on the observer’s celestial screen. However, Eq.(22) proves that in Modified Gravity (MOG), the massless scalar field wave directly modulates the critical impact parameter of the trapped null geodesics.
An observer monitoring this black hole with highly sensitive, next-generation interferometry would see the entire circular shadow rhythmically dilate and contract. To quantify this observable, we calculate the time-dependent fluctuation in the shadow’s total apparent area, $\delta A(t)$. By substituting the geometric expansion of the unperturbed squared shadow radius into the standard area formula $A=\pi R_{\text{sh}}^{2}$, we obtain the exact phenomenological expression for the area fluctuation:

### v2

# arXiv:2602.21432v2[gr-qc] 03 Mar 2026

@@ Abstract
In this paper, we investigate the dynamic phenomenological signatures of a Schwarzschild-MOG black hole shadow perturbed by passing gravitational waves. By perturbing the Hamilton-Jacobi equation for photon null geodesics, we demonstrate that the unique field content of MOG breaks the observational degeneracy with standard General Relativity. We mathematically prove two distinct, time-dependent signatures. First, the massless MOG scalar field induces a volumetric breathing-mode polarization, causing the total apparent area of the shadow to expand and contract rhythmically. Second, the massive MOG vector field undergoes quantum vacuum dispersion, arriving at the observer with a predictable time delay. The delayed massive-wave sources secondary longitudinal metric perturbations manifest as a sudden, asymmetric translational wobble of the shadow on the celestial screen. These dynamic geometric shifts offer a robust observational template for next-generation interferometry to strictly test the existence of massive force carriers and scalar fields in gravity.
@@ I Introduction
Among the most compelling alternative frameworks is Scalar-Tensor-Vector Gravity (STVG), widely referred to as Modified Gravity (MOG) [12]. In MOG, the gravitational interaction is not governed exclusively by a massless tensor field. Instead, the theory introduces an enhanced dynamical scalar gravitational coupling, together with a massive vector field that mediates a repulsive fifth force. This theoretical architecture successfully replicates the flat rotation curves of galaxies and the dynamics of galaxy clusters without necessitating cold dark matter [13, 14]. In the strong-field regime, the static MOG black hole solutions yield event horizons and photon spheres whose geometries are distinctly altered by the MOG deformation parameter $\alpha$, producing static observer shadows that differ slightly from their Schwarzschild and Kerr counterparts [15, 16, 17].
Although the static shadows of modified gravity theories have been extensively categorized, the universe is inherently dynamic [18]. A fundamental gap remains in understanding how these modified geometries respond to transient gravitational radiation. When a standard General Relativity gravitational wave washes over a black hole, it perturbs the spacetime metric exclusively through transverse-traceless tensor polarization modes [19]. These modes are strictly volume preserving. Therefore, if a standard gravitational wave dynamically distorts a black hole shadow, it can only stretch and squash the circular boundary into an ellipse, strictly conserving the total apparent area of the shadow on the observer’s celestial screen.
In this paper, we bridge this theoretical gap by mathematically deriving the dynamic interaction between gravitational waves and the shadow of a non-rotating Schwarzschild-MOG black hole. Because MOG introduces extra degrees of freedom into the gravitational sector, its gravitational waves carry novel polarization modes [20, 21]. We utilize the Hamilton-Jacobi formalism for perturbed null geodesics [22, 23] to explicitly map how the MOG scalar and massive vector fields dynamically shift the critical impact parameters of trapped photons. We demonstrate that the scalar field fluctuations generate a strictly volumetric breathing mode, meaning the area of the shadow will rhythmically dilate and contract [24, 25], a smoking gun signature forbidden in standard General Relativity.
Furthermore, we explore the phenomenological consequences of the massive vector field. Because massive fields undergo vacuum dispersion [26, 27], they travel at a group velocity strictly less than the speed of light. We show that this dispersion generates a delayed geometric echo. When this massive, delayed wave interacts with the black hole geometry, it excites secondary longitudinal metric perturbations, leading to an asymmetric translational wobble at the shadow’s center.
@@ II The Unperturbed Schwarzschild-MOG Spacetime
To understand how a gravitational wave dynamically alters a black hole shadow, we must first rigorously define the static, unperturbed background. In this section, we construct the spherically symmetric, non-rotating black hole spacetime in STVG, commonly referred to as MO) [12]. We then determine the exact radius of the photon sphere and the resulting static shadow projected to a distant observer [15].
In standard General Relativity, a non-rotating black hole is described by the Schwarzschild metric. In MOG, the presence of a massive vector field and an enhanced scalar gravitational constant alters the geometry [12]. The unperturbed Schwarzschild-MOG metric in standard spherical coordinates $(t,r,\theta,\phi)$ is given by the line element:
@@
The boundary of a black hole’s shadow is defined by the paths of photons that barely escape the black hole’s gravity. To trace these paths, we use the Hamilton-Jacobi equation for a photon traveling along a null geodesic, $ds^{2}=0$.
Because the spacetime is spherically symmetric, without loss of generality, confine the photon’s orbit to the equatorial plane by setting $\theta=\pi/2$. This simplifies the metric symmetries, yielding two fundamentally conserved quantities for the photon along its trajectory:
@@
The edge of the black hole shadow is formed by photons trapped in an unstable circular orbit called the photon sphere. For a photon to remain in a perfect circular orbit at a constant radius $r_{p}$, it must sit precisely at the peak of the effective potential barrier. This requires two simultaneous physical conditions:
@@
Applying the second condition by maximizing Eq. (8) yields a simple quadratic equation governing the photon sphere radius,
@@
A distant observer does not see $r_{p}$ directly. Instead, they see the apparent shadow projected onto their celestial sky [28]. The size of this dark circular disk is determined by the critical impact parameter, denoted as $b_{c}=L/E$. By rearranging the first condition, we find that the static shadow radius $R_{\text{sh}}$ is directly tied to the impact parameter:
@@
Substituting our MOG metric function $f(r_{p})$, the final unperturbed shadow radius becomes,
@@
Figure 1: Simultaneous evolution of the black hole shadow radius $R_{\text{sh}}$ (left axis) and the total apparent area fluctuation $\delta A(t)$ (right axis) under the influence of a transient scalar gravitational wave. The navy dashed curves represent the Schwarzschild (GR) baseline, while the crimson and green solid curves depict the enhanced Schwarzschild-MOG response for a deformation parameter $\alpha=0.2$. Both models exhibit a characteristic breathing mode triggered at the wave arrival time $t_{0}=5.0$, characterized by an exponentially damped oscillation ($\tau=8.0$) at the quasi-normal mode frequency $\omega_{b}=1.5$. Note that the MOG framework yields a larger static baseline and a greater amplitude of area fluctuations, providing a distinct observational signature of scalar-tensor-vector gravity in the strong-field regime.
In Fig. 1, we present the numerical results of the combined dynamics of the shadow radius and area fluctuations. The plot highlights the fundamental coupling between the scalar gravitational strain $h_{b}(t)$ and the geometric observables. Using a dual-axis representation, we show that the MOG deformation parameter $\alpha$ acts as a scaling factor for the breathing mode, not only is the static shadow larger than the Schwarzschild counterpart, but the magnitude of the dynamic response to the same incident wave is also amplified. This synchronization of the radial oscillation and the volumetric area change, mathematically derived in Eq. (22) and Eq.(26), constitutes a primary differentiator for the STVG framework.
@@ III Scalar Gravitational Wave Perturbations: The Breathing Mode
Having established the static Schwarzschild-MOG background, we now introduce a time-dependent perturbation to the spacetime. In standard General Relativity, gravitational waves possess only two transverse-traceless tensor polarization modes, known as the plus and cross modes. These tensor modes are strictly volume-preserving; if they pass through a circular ring of test particles, they will stretch and squash the circle into an ellipse, but the total area enclosed by the ring remains constant. However, MOG introduces an additional dynamic scalar field, $G$, into the gravitational action [29]. Fluctuations in this scalar field generate a unique scalar gravitational wave polarization known as the breathing mode [30, 31]. Unlike tensor modes, the breathing mode is entirely volumetric—it causes space to expand and contract uniformly in the transverse plane rhythmically.
@@
where $\bar{g}_{\mu\nu}$ is our static Schwarzschild-MOG background from Eq. (1), and $h_{\mu\nu}$ is the significantly small perturbation caused by the passing wave. For a pure scalar breathing mode generated by the MOG scalar field, the spatial perturbations in the transverse plane, the $x$-$y$ plane, take the form of a simple conformal factor. We denote this time-dependent strain as $h_{b}(t)$.
To analytically model the transient scalar strain $h_{b}(t)$, we draw upon the foundational principles of black hole perturbation theory. When a black hole is dynamically perturbed, such as during the final stages of a compact binary coalescence, the background spacetime geometry acts as an effective potential barrier to incoming wave fields. As first demonstrated by Vishveshwara [32] and generalized by Teukolsky [33], the newly formed or perturbed black hole radiates away these geometric deformations as QNMs. Because the black hole is an open, dissipative system losing energy to both spatial infinity and the event horizon, the characteristic eigenvalues governing these wave equations are strictly complex, taking the form $\tilde{\omega}=\omega_{b}-i/\tau$. Here, the real component $\omega_{b}$ dictates the fundamental oscillation frequency, while the imaginary component yields a characteristic exponential damping time $\tau$.
Translating these complex frequencies into the time domain yields a strictly damped sinusoid. Therefore, to model the incident scalar wave packet as a fundamental QNM ringdown, we define the time-dependent scalar strain $h_{b}(t)$ as a piecewise function. Applying the strict causality constraint of the wave’s arrival time $t_{0}$, the strain is explicitly:
@@
To determine how this oscillating metric affects the path of the light rays. The motion of a photon is dictated by the Hamiltonian,
@@
where $p_{\mu}$ is the four-momentum of the photon. We split this Hamiltonian into two parts: the unperturbed background part ($\bar{H}$) and the small shift caused by the wave ($\delta H$). It is expressed as $H=\bar{H}+\delta H=0$. The unperturbed Hamiltonian yields the static orbits we derived in Section II. The perturbation term, driven entirely by the gravitational wave, is:
@@
Because the wave is traveling along the $z$-axis, it primarily disrupts the angular momentum components of the photon’s momentum, specifically $p_{x}$ and $p_{y}$ in Cartesian coordinates. Substituting our perturbed inverse metric from Eq. (15) into the Hamiltonian perturbation, and recalling that the square of the transverse momentum corresponds to $L^{2}/r^{2}$ in spherical coordinates, yiels,
@@
The shift in the Hamiltonian directly modifies the radial equation of motion for the photon. When we substitute the perturbed Hamiltonian back into Eq. (7), the gravitational wave essentially injects a time-dependent fluctuation into the effective potential barrier. The new, total dynamic effective potential is the sum of the unperturbed potential (Equation 8) and the wave perturbation:
@@
Eq. (19) reveals a profound physical insight. Because the scalar wave strain $h_{b}(t)$ depends only on time and not on the radial coordinate $r$, the spatial derivative of the effective potential ($\partial_{r}V_{\text{eff}}^{\text{total}}$) is not shifted to first order. Consequently, the physical location of the photon sphere remains perfectly anchored at the static radius $r_{p}$ as shown in Eq. (10). The wave washes over the black hole, but the spherical boundary where light is trapped does not change its radial coordinate. However, the energy required for a photon to escape that trapped orbit is dynamically altered by the wave.
To find the new edge of the black hole shadow, we must recalculate the critical impact parameter. We let the peak of the new total effective potential equal to the squared energy of the photon ($E^{2}$) evaluated at the exact photon sphere radius $r_{p}$:
@@
It is critical to clarify the physical regime in which this formulation operates explicitly. By setting the peak of the time-dependent effective potential equal to the unperturbed squared energy $E^{2}$ in Eq. (20), our derivation implicitly relies on the adiabatic approximation. In a strictly dynamic spacetime where the metric fluctuates with time ($\partial_{t}g_{\mu\nu}\neq 0$), the photon’s total energy $E=-p_{t}$ is not perfectly conserved along its null geodesic. This adiabatic assumption is most robust when the period of the passing gravitational wave is significantly longer than the time it takes for a photon to cross the strong-field region of the photon sphere. We acknowledge that for highly dynamic perturbations such as fundamental quasinormal modes where the wave’s oscillation frequency ($\omega_{b}\sim 1/M$) is comparable to the photon’s orbital frequency the photon can exchange energy with the gravitational wave, meaning $dE/d\lambda\neq 0$. In such extreme regimes, achieving mathematical exactness would require a fully integrated geodesic equation. Nevertheless, evaluating the critical impact parameter through this adiabatic Hamiltonian formalism successfully and elegantly captures the leading-order kinematic deformation of the shadow’s boundary, providing a highly accurate phenomenological baseline. By rearranging this expression to solve for the time-dependent critical impact parameter, $b_{c}(t)=L/E$, we obtain the mathematical formula for the dynamic shadow radius, $R_{\text{sh}}(t)$:
@@
Because the amplitude of a gravitational wave strain is significantly small, $h_{b}\ll 1$), we can safely apply a first-order Taylor expansion to the fraction $(1-h_{b})^{-1}\approx 1+h_{b}(t)$. Taking the square root of both sides, and recognizing that $r_{p}/\sqrt{f(r_{p})}$ is exactly the unperturbed static shadow radius $\bar{R}_{\text{sh}}$ from Eq. (12), we yield a dynamics shadow radius,
@@
To provide a observational testing, we synthesize the static modified geometry with our dynamic wave model. By substituting the geometric expansion of the unperturbed shadow radius and the quasinormal mode strain $h_{b}(t)$ into Eq. (22), we arrive at the master equation for the dynamic shadow radius,
@@
This expanded expression elegantly decouples the theoretical background from the transient perturbation. The leading square-root term isolates the static Schwarzschild-MOG geometry, explicitly governed by the Newtonian mass $M$, the deformation parameter $\alpha$, and the critical photon sphere $r_{p}$. To fully unpack the localized gravitational wave interaction, it is most instructive to evaluate Eq. (23) across its two distinct temporal regimes. Let us explicitly define the unperturbed, static Schwarzschild-MOG shadow radius as the geometric baseline,
@@
In standard General Relativity, a passing gravitational wave perturbs the spacetime metric exclusively through transverse-traceless tensor modes. Consequently, such a wave can only dynamically distort the black hole shadow into an ellipse, strictly conserving its total apparent area on the observer’s celestial screen. However, Eq.(22) proves that in MOG, the massless scalar field wave directly modulates the critical impact parameter of the trapped null geodesics.
An observer monitoring this black hole with highly sensitive next-generation interferometry would observe the entire circular shadow dilate and contract in rhythm. To quantify this observable, we calculate the time-dependent fluctuation in the shadow’s total apparent area, $\delta A(t)$. By substituting the geometric expansion of the unperturbed squared shadow radius into the standard area formula $A=\pi R_{\text{sh}}^{2}$, we obtain the area fluctuation equation given as,